OpenStudy (anonymous):

18 is what percent of 27? how do i do the proportion

OpenStudy (unklerhaukus):

\[{18\over27}\times{ 100\%}\]

OpenStudy (unklerhaukus):

\[=\frac{2}{3}\times 100\%\]

OpenStudy (anonymous):

where does the x go??

OpenStudy (anonymous):

so you just multiply accros

OpenStudy (anonymous):